two objects are directly in line with a cliff which stands 282ft. above the water. there angles of depression from the top of the cliff are 42 degree 25 minutes and 37 degress 15 minutes. how far apart are the objects and what are their horizontal distance from the cliff?

Answer by mananth(16949)


Let x be the distance between the two objects
a is the distance from the object nearer to the cliff
Tan 42 25' = (x+a)/282
Tan 42.42
Tan 37 15' = a/282
Tan 37.25


Tan42.42-Tan 37.25= (x+a)/282- a/282
0.91-0.76= x/282
0.15*282=x
x= 42.3
CHECK
a=214.3
(x+a)=256.6
Tan 42.42= (x+a)/282
256.6/282= 0.91
Tan 42.42 = 0.91
Hence proved
m.ananth@hotmail.ca